Determine the appropriate base of each logarithm
log base b (1/15)=-1.9
log base b 80=3.9
log base b 6=0.9
log base b 66=3.1

Answer by Alan3354(69443)

Determine the appropriate base of each logarithm
log base b (1/15)=-1.9
-----------
---> b^-1.9 = 1/15
Solve for b
Using ln, natural logs, base e:
-1.9*ln(b) = ln(1/15)
ln(b) = ln(1/15)/(-1.9)
ln(b) = -2.708/(-1.9) = ~1.4252
b = e^(1.4252)
b =~ 4.159
--------------
===================
log base b 80=3.9
---
b^3.9 = 80
Using common logs, base 10:
3.9*log(b) = log(80)
log(b) = log(80)/3.9 = ~0.48797
b = 10^(0.48797) = ~3.0759
-------
log base b 6=0.9
log base b 66=3.1